OpenStudy (anonymous):

A survey found that 15 out of 100 people have blue eyes. On the basis of this survey, how many people in a group of 5000 are likely to have blue eyes? A) 75 B) 150 C) 750 D) 1500

OpenStudy (anonymous):

15/100 = x /5000 [Set up as a ratio, sorry can't figure out the equation device on here!] .15 = x/5000 .15(5000) = x 750 = x C is your answer.

OpenStudy (anonymous):

@Bribear thank you so much! your awesome my friend!!!! :)

OpenStudy (anonymous):

You're very welcome ^^